Some European trucks run on energy stored in a rotating flywheel, with an electric motor getting the flywheel up to its top speed of 180 πrad/s. One such flywheel is a solid, uniform cylinder with a mass of 310 kg and a radius of 1.05 m. (a) What is the kinetic energy of the flywheel after charging? (b) If the truck uses an average power of 9.4 kW, for how many minutes can it operate between chargings?
Answer #1

here,

the top speed , w = 180 pi rad/s

mass ,m = 310 kg , radius , r = 1.05 m

a)

the kinetic energy of the flywheel after charging , KE = 0.5 * I * w^2

KE = 0.5 * 0.5 * 310 * 1.05^2 * (180 pi)^2 J

KE = 2.73 * 10^7 J

b)

average power , P = 9.4 KW = 9400 W

let the time taken be t

t = KE/P

t = 2.73 * 10^7 /( 9400) s

t = 2903.7 s

t = 48.4 min

the time for which it can be operted between charging is 48.4 min